OpenStudy (anonymous):

integral from 0 to 3 5x Sin (3x+1) I'm pretty sure this is an integration by parts. I made U 5x and dv = sin (3x+1). Is this the correct path?

OpenStudy (astrophysics):

\[\int\limits_{0}^{3} 5x \sin(3x+1)dx\] like this?

OpenStudy (welshfella):

yes by parts would do it

OpenStudy (astrophysics):

By parts should work, u = 5x, dv = sin(3x+1) so now integrate for v

OpenStudy (astrophysics):

Note: \[\int\limits u dv = uv - \int\limits v du \]

OpenStudy (anonymous):

so would v end up being \[\frac{ 1 }{ 3 } \sin (3x+1) \] sorry for the dumb question. Been a couple of years since Calculus 1!

OpenStudy (anonymous):

oops, change that sin to a -cos.

OpenStudy (astrophysics):

Yup

OpenStudy (astrophysics):

-1/3cos(3x+1)

OpenStudy (astrophysics):

Keep going, I'm sure you'll get it!